Calculate the number of ATP produced from the complete oxidation of myristic acid.

Each round of the citric acid cycle uses one molecule of acetyl-CoA. How many molecules of ATP are produced from all of the acetyl-CoA formed from myristic acid going through the citric acid cycle?
Answer #1

Fatty acid oxidation takes place mostly is beta oxidation. Myristic acid is a 14 C compound which undergoes many rounds of cycles until it attains 2c compound acetyl coA. therefore it undergoes 6 cycles and yields 7 molecules of acetyl coA. In each cycle one NADH and one FADH2 is produced. One NADH yields 3 ATP and one FADH2 yields 2 ATP on entering electron transport chain. On the whole 6 NADH and 6 FADH2 are produced from complete oxidation of myristic acid.  

No. Of ATP molecules generated = (6 x 3) + (6x2)

= 18+ 12

=30

2 Atp molecules are utilized for fatty acid oxidation activation.

Therefore 30 - 2 = 28 ATP molecules are generated.

Krebs cycle yields 3 NADH and 1 FADH2 and 1 ATP OR GTP on utilizing one acetyl Co A.

7 acetyl coA yields 21 NADH and 7 FADH2 and 7 ATP

Therefore total no.of ATP yield = (3 x 21 ) + ( 7 x 2) + 7

= 63 + 14 +7

= 84 ATP molecules are produced

The complete oxidation of myristic acid yields 30 ATP molecules. On utilizing 2 ATP for activation, the remaining ATP molecules are 28 ATP.

From the 7 acetyl CoA generated from myristic acid oxidation, 84 ATP molecules are generated.